# Monotonic Path-Specific Effects: Application to Estimating Educational Returns

### v1

@@ 2.2 Generalization to $K$ Intermediate Transitions
where $\Delta_{K}=\tau_{K}$, i.e. $\Delta_{K}$ is a gross or continuation effect, since this latter path is a composite one that contains all residual paths omitted in the decomposition (i.e., through educational transitions subsequent to $K$, if they exist). Thus, the $\theta_{k}$ terms capture how much of the total effect of high school completion flows through each intermediate transition considered (i.e., via college attendance, via college completion, and via graduate school attendance), while $\Delta_{0}$ captures that portion of the total effect that operates directly, net of the $K$ intermediate transitions considered.
@@ 2.3 Identification
To identify the causal effects of interest, I rely on a series of sequential ignorability assumptions. While most closely associated with the dynamic treatment effects literature, which rely on observing a complete set of time-varying confounders in order to identify longitudinal effects (see e.g. Lewis and Syrgkanis, 2020; Viviano and Bradic, 2021; Bodory et al., 2022), these assumptions can be transferred to a mediation context, given the fact that the mediators of interest are all causally ordered. As will be discussed in the following section, sequential ignorability identification assumptions are distinct from - and in fact weaker than - the assumptions typically employed in studies of causal mediation.
Before proceeding, I introduce the following shorthands. Let $M_{0}\triangleq A$ and $M_{k}=\emptyset\forall k<0$. In order to estimate the decomposition shown in Equation 5, it suffices to identify the expectation of two types of composite counterfactuals ($Y(\overline{1}_{k},m_{k+1})$ and $M_{k+1}(\overline{1}_{k+1})$), as well as covariance terms of the form $\mathrm{cov}[M_{k+1}(\overline{1}_{k+1}),Y(\overline{1}_{k+2})-Y(\overline{1}_{k+1},0)]$ $\forall k\in[K-1]$. I invoke the following three assumptions:
@@
Consistency: for any unit, if $A=a$, $Y=Y(a)$; if $(A,\overline{M}_{k})=\{\overline{1}_{k},m_{k}\}$, then $Y=Y(\overline{1}_{k},m_{k})$ $\forall k\in[K]$, and if $(A,\overline{M}_{k})=\overline{1}_{k+1}$, then $M_{k+1}=M_{k+1}(\overline{1}_{k+1})\forall m_{k+1}\in\{0,1\},\forall k\in[K-1]$.
@@
Sequential ignorability: $(M_{1}(1),Y(a))\perp\!\!\!\perp A|X$; $Y(\overline{1}_{k},m_{k})\perp\!\!\!\perp\overline{M}_{k}|X,\overline{Z}_{k},\overline{M}_{k-1}$ and $M_{k+1}(\overline{1}_{k+1})\perp\!\!\!\perp\overline{M}_{k}|X,\overline{Z}_{k},\overline{M}_{k-1},\forall m_{k}\in\{0,1\},\forall k\in\{1,\dots,K\},M_{0}\equiv A$.
@@
Positivity: $p_{A|X}(a|x)>\epsilon>0$, $p_{M_{k}|X,A,\overline{Z}_{k},\overline{M}_{k-1}}(m_{k}|x,a,\overline{z}_{k},\overline{m}_{k-1})>\epsilon>0$ $\forall k\in[K]$.
Assumption 3 (consistency) states that a unit’s observed outcome equals its potential outcome under a given treatment sequence. Note that under under the Assumption 1 (Composition), if $Y=Y(\overline{1}_{k},m_{k})$, then $Y=Y(\overline{1}_{k},m_{k},M_{k+1}(\overline{1}_{k},m_{k}))=Y(\overline{1}_{k},m_{k},M_{k+1}(\overline{1}_{k},m_{k})),$ $\dots,M_{K}(\overline{1}_{k},m_{k},M_{k+1}(\overline{1}_{k},m_{k}),\dots M_{K-1}((\overline{1}_{k},m_{k}),M_{k+1}(\overline{1}_{k},m_{k}),\dots,M_{k-2}(\dots))$.
In plain words, the KkK-k mediators after mediator $k$ all take their natural levels. Assumption 4 (sequential ignorability) is the no unmeasured confounding assumption for the treatment and all mediators. It is considered plausible when sufficient pre-treatment and intermediate covariates $(X,\overline{Z}_{K})$ are collected. Finally, Assumption 5 (positivity) requires that treatment and mediator assignment is not deterministic. Under Assumptions 3-5, $\mathbb{E}[Y(\overline{1}_{k},m_{k})]$ and $\mathbb{E}[M_{k+1}(\overline{1}_{k+1})]$ are identified, respectively, as
$\mathbb{E}[Y(\overline{1}_{k},m_{k})]=\int_{x}\int_{\overline{z}_{k}}\mathbb{E}[Y|x,\overline{z}_{k},\overline{1}_{k},m_{k}]\big{[}\prod_{j=1}^{k}dP(z_{j}|x,\overline{z}_{j-1},\overline{1}_{j-1})\big{]}dP(x)$ (6)
$\mathbb{E}[M_{k+1}(\overline{1}_{k+1})]=\int_{x}\int_{\overline{z}_{k}}\mathbb{E}[M_{k+1}|x,\overline{z}_{K},\overline{1}_{k+1}]\big{[}\prod_{j=1}^{k}dP(z_{j}|x,\overline{z}_{j-1},\overline{1}_{j-1})\big{]}dP(x)$ (7)
@@
$\eta_{k}=\tau_{k-1}-\Delta_{k-1}-\pi_{k}\tau_{k}.$
The above decomposition has an analog in the context of a mediation–based decomposition of the ATE with multiple ordered mediators, but differs from conventional mediation analysis in important ways. To illustrate the differences, consider a binary treatment, $A$, an outcome of interest, $Y$, and a vector of pretreatment covariates, $X$, and let $M_{1},M_{2},\ldots M_{K}$ denote $K$ causally ordered mediators, assuming that for any $i<j,M_{i}$ precedes $M_{j}$, as above. Moreover, let an overbar denote a vector of variables, so that $\bar{M}_{k}=(M_{1},M_{2},\ldots M_{k})$$\bar{m}_{k}=(m_{1},m_{2},\ldots m_{k})$, and $\bar{a}_{k}=(a_{1},a_{2},\ldots a_{k})$. Using the potential outcomes notation as above, we can define the following expectation of a nested counterfactual,
$\psi_{a\bar{a}_{k}}\triangleq\mathbb{E}\left[Y(a,\bar{M}_{k}(\bar{a}_{k}))\right],$
where $\bar{M}_{k}(\bar{a}_{k})\triangleq\big{(}\bar{M}_{k-1}(\bar{a}_{k-1}),M_{k}(a_{k},\bar{M}_{k-1}(\bar{a}_{k-1}))\big{)},\forall k\in[K]$. Under Pearl’s (2009) nonparametric structural equation model (NPSEM), Zhou (2022b) demonstrates that the ATE of $A$ on $Y$ can be decomposed into $K+1$ identifiable path-specific effects (PSEs) corresponding to each of the causal paths $A\rightarrow Y$ and $A\rightarrow M_{k}\leadsto Y$ $(k\in[K])$:
$\mathrm{ATE}=\psi_{\overline{1}}-\psi_{\overline{0}}=\underbrace{\psi_{1,\overline{0}_{K},}-\psi_{\overline{0}_{K+1}}}_{A\rightarrow Y}+\sum_{k=1}^{K}\underbrace{\big{(}\psi_{\overline{1}_{k+1},\underline{0}_{k+1}}-\psi_{\overline{1}_{k},\underline{0}_{k+1}}}_{A\rightarrow M_{k}\rightsquigarrow Y}\big{)}.$ (8)
This decomposition holds algebraically when Assumption 2 does not hold (i.e., when the mediators are not monotonic). In contrast, the monotonic characteristic of the proposed decomposition leads to several important differences. First, the PSE decomposition of the ATE in general mediation settings is not algebraically unique, and thus the PSEs defined under alternative decompositions will differ if the effects of the treatment and each mediator vary across levels of the other mediators. In fact, depending on the order in which the paths $A\rightarrow Y$ and $A\rightarrow M_{k}\rightsquigarrow Y$ are considered, there are $(K+1)!$ identifiable different ways of decomposing the ATE; the decomposition shown in Equation 8 is just one such decomposition. Consider the case of two causally dependent mediators. In this setting, the causal pathway $A\to M_{2}\rightsquigarrow Y$ can be defined with respect to four different combinations of levels of the treatment and first mediator: under (i) $a=1$ and $M_{1}(1)$, (ii) $a=1$ and $M_{1}(0)$, (iii) $a=0$ and $M_{1}(1)$, or (iv) $a=0$ and $M_{1}(0)$. By contrast, as a direct consequence of monotonicity, the MPSE decomposition is the unique PSE decomposition of the ATE.
Second, for general PSE decompositions of the ATE, the set of identifiable decompositions is merely a small subset of the total number of decompositions that hold algebraically (see Avin et al., 2005). In particular, the identifiable decomposition does not enable us to disentangle the mediating effects of $M_{k}$ that are direct (net of subsequent mediators) and indirect (through different combinations of subsequent mediators). For example, in the case of two causally dependent mediators, to assess the mediating role of $M_{1}$, only the composite path $A\rightarrow M_{1}\rightsquigarrow Y=(A\rightarrow M_{1}\rightarrow Y)+(A\rightarrow M_{1}\rightarrow M_{2}\rightarrow Y)$ is identified. By contrast, mediator monotonicity permits a finer-grained decomposition of the ATE: each PSE is identified. In the case of two causally dependent mediators, for example, the causal path $A\rightarrow M_{2}\rightarrow Y$ is zero, and as a result, each of the paths $A\to Y$, $A\rightarrow M_{1}\rightarrow Y$ and $A\rightarrow M_{1}\rightarrow M_{2}\rightarrow Y$ are identifiable. Figure 1 illustrates the causal pathways defined and identified under the proposed decomposition in the case of two monotonic mediators.
Finally, the sequential ignorability assumption required to identify the MPSE decomposition is weaker than those required to identify a generic PSE decomposition of the ATE. Specifically, the latter requires Pearl’s (2009) non-parametric structural equation model (NPSEM), which stipulates that $\left(M_{k+1}\left(a_{k+1},\bar{m}_{k}\right),\ldots M_{K}\left(a_{K},\bar{m}_{K-1}\right),Y\left(a_{K+1},\bar{m}_{K}\right)\right)\perp\!\!\!\perp M_{k}\left(a_{k},\bar{m}_{k-1}^{*}\right)\mid X,A,\bar{M}_{k-1},\forall k\in[K].$ This assumption, sometimes referred to as the “cross-world” independence assumption, is stronger than the sequential ignorability assumption (4) required to identify the MPSE decomposition since it rules out the existence of confounders of the mediators, be they observed or unobserved. By contrast, the MPSE decomposition identification results accommodate observed intermediate confounding without altering the substance of the decomposition.
3 Semiparametric, EIF-Based Estimation
The identification results outlined above suggest that the proposed decomposition can be estimated via several approaches, including outcome-based modeling, models for the treatment and mediators via inverse probability weighting, as well as doubly robust approaches. Parametric procedures are attractive because of their conceptual simplicity and ease of implementation: in Supplementary Material A I show how, under a set of linear models for the outcome and mediators, the $\theta_{k}$ components in Equation 5 can be read off from simple functions of coefficients in these linear models. However when $X$ and $\overline{Z}_{K}$ are high-dimensional, parametric estimators which require a user-defined specification of the data-generating process may suffer biases resulting from model misspecification. In order to reduce model dependency, in this section I provide a nonparametric estimation approach which draws on a debiased machine learning (DML) approach. My DML approach is characterized by two components: first, the use of a Neyman orthogonal estimating equation based on the efficient influence function (EIF) for the target parameters, which makes estimates of the parameter “locally robust” to estimates of the nuisance functions; second, the use of a $K$-fold cross-fitting algorithm (Chernozhukov et al., 2017).

### v2

@@ 2.2 Generalization to $K$ Intermediate Transitions
where $\pi_{0}=1$. Further, $\Delta_{K}=\tau_{K}$ is a gross or continuation effect, since this latter path is a composite one that contains all residual paths omitted in the decomposition (i.e., through educational transitions subsequent to $K$, if they exist). Thus, the $\theta_{k}$ terms capture how much of the total effect of high school completion flows through each intermediate transition considered (i.e., via college attendance, via college completion, and via graduate school attendance), while $\Delta_{0}$ captures that portion of the total effect that operates directly, net of the $K$ intermediate transitions considered.
The decomposition introduced in the previous section has an analog in the context of a mediation–based decomposition of the ATE with multiple ordered mediators, but differs from these conventional mediation analyses in important ways. To illustrate the differences, consider a binary treatment, $A$, an outcome of interest, $Y$, and a vector of pretreatment covariates, $X$, and let $M_{1},M_{2},\ldots M_{K}$ denote $K$ causally ordered mediators, assuming that for any $i<j,M_{i}$ precedes $M_{j}$, as above. Moreover, define $M_{0}\equiv\emptyset$, and let
$\overline{M}_{k}\equiv(M_{1},M_{2},\ldots,M_{k}),\qquad\overline{M}_{k}(a)\equiv\big(M_{1}(a),M_{2}(a,M_{1}(a)),\ldots,M_{k}(a,\overline{M}_{k-1}(a))\big),$
with $\overline{M}_{0}(a)\equiv\emptyset$. Using the potential outcomes notation as above, define the following expectation of a nested counterfactual,
$\psi(a,\overline{M}_{k}(a^{\ast}))\triangleq\mathbb{E}\!\left[Y\!\big(a,\overline{M}_{k}(a^{\ast})\big)\right].$
Under Pearl’s (2009) nonparametric structural equation model (NPSEM), Zhou (2022b) demonstrates that the ATE of $A$ on $Y$ can be decomposed into $K+1$ identifiable path-specific effects (PSEs) corresponding to each of the causal paths $A\rightarrow Y$ and $A\rightarrow M_{k}\rightsquigarrow Y$ $(k\in[K])$ (see also Daniel et al., 2015):
$\mathrm{ATE}=\mathbb{E}[Y(1)-Y(0)]=\underbrace{\psi(1,\overline{M}_{K}(0))-\psi(0,\overline{M}_{K}(0))}_{A\rightarrow Y}+\sum_{k=1}^{K}\underbrace{\big(\psi(1,\overline{M}_{k}(0))-\psi(1,\overline{M}_{k-1}(0))\big)}_{A\rightarrow M_{k}\rightsquigarrow Y}.$ (7)
This decomposition holds algebraically when Assumption 2 does not hold (i.e., when the mediators are not monotonic), and differs from the proposed decomposition (Equation 6) in several ways. [^4] Vansteelandt and Daniel (2017) propose a decomposition of the total effect in settings with multiple, potentially causally ordered mediators using interventional (in)direct effects. A key advantage of this approach is that, like the MPSE decomposition, it permits exposure-induced intermediate confounding by defining effects in terms of stochastic interventions on mediator distributions. When mediators are causally dependent, however, mediation through a given mediator is not, in general, fully summarized by the corresponding interventional indirect effect alone; an additional component capturing exposure-induced changes in the dependence between mediators (their Equation (8)) is required to complete the decomposition. Accordingly, in settings with causally dependent mediators, mediation along a particular sequential path may be zero even when the associated interventional indirect effect is nonzero. By contrast, the framework proposed here exploits monotonicity to restrict the support of downstream mediators to that of upstream transitions, so that all mediation operates along well-defined sequential paths and no separate dependence term is required. Interestingly, under monotonicity, the interventional direct effect in Vansteelandt and Daniel (2017) coincides exactly with $\Delta_{0}$ in the MPSE decomposition, echoing the result that the randomized natural direct effect at zero equals $\Delta_{0}$, as discussed in Section 2.1.
First and most importantly, the monotonic decomposition departs from conventional mediation decompositions in distinguishing a different set of causal pathways. As shown in the previous section, the monotonic and standard mediation decompositions both begin with a natural effect decomposition in the case of a single mediator. Conventional mediation decompositions then obtain a 2-mediator decomposition by decomposing the direct path $A\to Y$ net of $M_{1}$ into its direct component net of $M_{2}$ and indirect path $A\to M_{2}\to Y$. This is shown in Figure 1, Panel A. Notably, to assess the mediating role of $M_{1}$, only the composite path $A\rightarrow M_{1}\rightsquigarrow Y=(A\rightarrow M_{1}\rightarrow Y)+(A\rightarrow M_{1}\rightarrow M_{2}\rightarrow Y)$ is identified. The reason for this is that identification of the pure path-specific effects $A\rightarrow M_{1}\rightarrow Y$ or $A\rightarrow M_{1}\rightarrow M_{2}\rightarrow Y$ fails when the recanting witness criterion is violated (Avin et al., 2005). Specifically, $M_{1}$ acts as a “recanting witness” because it lies on the path $A\rightarrow M_{1}\rightarrow Y$ but also has an additional path to $Y$ through $M_{2}$ ($M_{1}\rightarrow M_{2}\rightarrow Y$) that is not contained in the path of interest. The nested counterfactual $\mathbb{E}[Y(a,M_{1}(a_{1}),M_{2}(a_{2},M_{1}(a_{12})))]$ is identified if and only if $a_{1}=a_{12}$. Consequently, the individual PSEs for $A\rightarrow M_{1}\rightarrow Y$ and $A\rightarrow M_{1}\rightarrow M_{2}\rightarrow Y$ are not identified, whereas the composite path $A\rightarrow M_{1}\rightsquigarrow Y$, which includes all possible paths from $M_{1}$ to $Y$, remains identifiable. More generally, traditional decompositions do not disentangle the mediating effects of $M_{k}$ that are direct (net of subsequent mediators) and indirect (through different combinations of subsequent mediators).
By contrast, as shown in Figure 1, Panel B, my proposed decomposition under mediator monotonicity begins with the natural effect decomposition with a single mediator as in Equation 1, and then decomposes the path via $M_{1}$ into pathways that operate further via $M_{2}$ ($A\rightarrow M_{1}\rightarrow M_{2}\rightarrow Y$) and that operate directly to $Y$ ($A\rightarrow M_{1}\rightarrow Y$), leaving the direct effect ($\Delta_{0}$) untouched. In fact, $\Delta_{0}$ is not further decomposable under monotonicity because this assumption yields a more restricted set of causal paths. To see why, consider the pathway $A\to M_{2}\to Y$ under monotonicity in the two-mediator setting:
$\mathbb{E}\!\left[Y\!\big(1,\,M_{1}(0),\,M_{2}(1,M_{1}(0))\big)-Y\!\big(1,\,M_{1}(0),\,M_{2}(0,M_{1}(0))\big)\right].$
Under mediator monotonicity, $M_{1}(0)=0$ (e.g., without completing high school one cannot attend college), and $M_{2}(\cdot,0)=0$ (e.g., without college attendance one cannot complete a BA). Therefore
$M_{2}(0,M_{1}(0))=M_{1}(0,0)=0.$
and the path $A\to M_{2}\to Y$ equals
Y(1,0,0)Y(1,0,0)=0.Y(1,0,0)-Y(1,0,0)=0.
More generally, consider the path-specific effect attributed to the pathway $A\to M_{k}\to Y$ with $k\geq 2$, defined in the usual way by fixing all earlier mediators at their natural levels under $A=1$ and varying only $M_{k}$ with respect to $A$:
$\tau_{A\to M_{k}\to Y}(a)=\mathbb{E}\!\left[Y\!\big(1,\overline{M}_{k-1}(0),M_{k}(1,\overline{M}_{k-1}(0))\big)-Y\!\big(1,\overline{M}_{k-1}(0),M_{k}(0,\overline{M}_{k-1}(0))\big)\right],$
where $\overline{M}_{k}(a)$ denotes the vector $(M_{1}(a),\ldots,M_{k}(a,\overline{M}_{k-1}(a)))$. Under monotonicity, fixing earlier mediators at their baseline levels implies that $\overline{M}_{k}(0)=0$, and that $M_{k}(1,\overline{M}_{k-1}(0))=M_{k}(0,\overline{M}_{k-1}(0))=0$. In words, there is no direct effect of the treatment on later mediators once earlier transitions are held fixed at their natural value under $A=0$. As a result, the two potential outcomes inside the expectation coincide, and
$\tau_{A\to M_{k}\to Y}(1)=0\qquad\text{for all }k\geq 2.$
Any indirect effect involving $M_{k}$ must therefore operate through the full causal chain $A\to M_{1}\to\cdots\to M_{k}\to Y$. Figure 2 illustrates the causal pathways defined and identified under the proposed decomposition in the case of two monotonic mediators.
Second, the PSE decomposition of the ATE in general mediation settings is not algebraically unique, and thus the PSEs defined under alternative decompositions will differ if the effects of the treatment and each mediator vary across levels of the other mediators. In fact, depending on the order in which the paths $A\rightarrow Y$ and $A\rightarrow M_{k}\rightsquigarrow Y$ are considered, there are $(K+1)!$ identifiable different ways of decomposing the ATE; the decomposition shown in Equation 7 is just one such decomposition. Consider the case of two causally dependent mediators. In this setting, the causal pathway $A\to M_{2}\rightsquigarrow Y$ can be defined with respect to four different combinations of levels of the treatment and first mediator: under (i) $a=1$ and $M_{1}(1)$, (ii) $a=1$ and $M_{1}(0)$, (iii) $a=0$ and $M_{1}(1)$, or (iv) $a=0$ and $M_{1}(0)$. By contrast, as a direct consequence of monotonicity, the MPSE decomposition is the unique PSE decomposition of the ATE.
Finally, the sequential ignorability assumption required to identify the MPSE decomposition is weaker than those required to identify a generic PSE decomposition of the ATE. Specifically, the latter requires Pearl’s (2009) non-parametric structural equation model (NPSEM), which stipulates that $\left(M_{k+1}\left(a_{k+1},\bar{m}_{k}\right),\ldots M_{K}\left(a_{K},\bar{m}_{K-1}\right),Y\left(a_{K+1},\bar{m}_{K}\right)\right)\perp\!\!\!\perp M_{k}\left(a_{k},\bar{m}_{k-1}^{*}\right)\mid X,A,\bar{M}_{k-1},\forall k\in[K].$ This assumption, sometimes referred to as the “cross-world” independence assumption, is stronger than the sequential ignorability assumption (4) required to identify the MPSE decomposition since it rules out the existence of confounders of the mediators, be they observed or unobserved. By contrast, the MPSE decomposition identification results accommodate observed intermediate confounding without altering the substance of the decomposition. I discuss this point in detail in the following section.
@@ 2.4 Identification
To identify the causal effects of interest, I rely on a series of sequential ignorability assumptions. While most closely associated with the dynamic treatment effects literature, which rely on observing a complete set of time-varying confounders in order to identify longitudinal effects (see e.g. Lewis and Syrgkanis, 2020; Viviano and Bradic, 2021; Bodory et al., 2022), these assumptions can be transferred to a mediation context, given the fact that the mediators of interest are all causally ordered. As discussed in the previous section, sequential ignorability identification assumptions are distinct from – and in fact weaker than – the assumptions typically employed in studies of causal mediation.
Let $M_{k}=\emptyset\forall k<1$. In order to estimate the decomposition shown in Equation 6, it suffices to identify the expectation of two types of composite counterfactuals ($Y(1,\overline{1}_{k-1},m_{k})$ and $M_{k+1}(1,\overline{1}_{k})$), as well as covariance terms of the form $\mathrm{cov}[M_{k+1}(1,\overline{1}_{k}),Y(1,\overline{1}_{k+1})-Y(1,\overline{1}_{k},0)]$ $\forall k\in[K-1]$. I invoke the following three assumptions:
@@
Consistency: for any unit, if $A=a$, $Y=Y(a)$; if $(A,\overline{M}_{k-1},M_{k})=(1,\overline{1}_{k-1},m_{k})$, then $Y=Y(1,\overline{1}_{k-1},m_{k})$ $\forall k\in[K]$, and if $(A,\overline{M}_{k})=(1,\overline{1}_{k})$, then $M_{k+1}=M_{k+1}(1,\overline{1}_{k})\forall m_{k+1}\in\{0,1\},\forall k\in[K-1]$.
@@
Sequential ignorability: $(M_{1}(1),Y(a))\perp\!\!\!\perp A|X$;
$Y(1,\overline{1}_{k-1},m_{k})\perp\!\!\!\perp{M}_{k}|X,\overline{Z}_{k},\overline{M}_{k-1}=\overline{1}_{k-1}$ and $M_{k+1}(1,\overline{1}_{k})\perp\!\!\!\perp{M}_{k}|X,\overline{Z}_{k},\overline{M}_{k-1}=\overline{1}_{k-1}$, $\forall m_{k}\in\{0,1\},\forall k\in\{1,\dots,K\},M_{0}\equiv A$.
@@
Positivity: $p_{A|X}(a|x)>\epsilon>0$, $p_{M_{k}|X,A,\overline{Z}_{k},\overline{M}_{k-1}}(m_{k}|x,a,\overline{z}_{k},\overline{m}_{k-1}=\overline{1}_{k-1})>\epsilon>0$ $\forall k\in[K]$.
Assumption 3 (consistency) states that a unit’s observed outcome equals its potential outcome under a given treatment sequence. Note that under Assumption 1 (Composition), if $Y=Y(1,\overline{1}_{k-1},m_{k})$, then $Y=Y(1,\overline{1}_{k-1},m_{k},M_{k+1}(1,\overline{1}_{k-1},m_{k}))=\dots=Y(1,\overline{1}_{k-1},m_{k},M_{k+1}(1,\overline{1}_{k-1},m_{k}),\dots,M_{K}(\cdot))$. In plain words, the KkK-k mediators after mediator $k$ all take their natural levels. Assumption 4 (sequential ignorability) is the no unmeasured confounding assumption for the treatment and all mediators. It is considered plausible when sufficient pre-treatment and intermediate covariates $(X,\overline{Z}_{K})$ are collected. Finally, Assumption 5 (positivity) requires that treatment assignment is not deterministic, and that mediator assignment is not deterministic when the treatment and prior mediators $\overline{m}_{k-1}$ take on a value of $1$.
Under Assumptions 35, $\mathbb{E}[Y(1,\overline{1}_{k-1},m_{k})]$ and $\mathbb{E}[M_{k+1}(1,\overline{1}_{k})]$ are identified, respectively, as
$\mathbb{E}[Y(1,\overline{1}_{k-1},m_{k})]=\int_{x}\int_{\overline{z}_{k}}\mathbb{E}[Y|x,\overline{z}_{k},1,\overline{1}_{k-1},m_{k}]\big[\prod_{j=1}^{k}dP(z_{j}|x,\overline{z}_{j-1},\overline{1}_{j-1})\big]dP(x)$ (8)
$\mathbb{E}[M_{k+1}(1,\overline{1}_{k})]=\int_{x}\int_{\overline{z}_{k}}\mathbb{E}[M_{k+1}|x,\overline{z}_{K},1,\overline{1}_{k}]\big[\prod_{j=1}^{k}dP(z_{j}|x,\overline{z}_{j-1},\overline{1}_{j-1})\big]dP(x)$ (9)
@@
$\eta_{k}=\tau_{k-1}-\Delta_{k-1}-\pi_{k}\tau_{k}.$ (10)
3 Estimation
The identification results outlined above suggest that the proposed decomposition can be estimated via several approaches, including outcome-based modeling, models for the treatment and mediators via inverse probability weighting, as well as doubly robust approaches. This section outlines two complementary estimation strategies (one semiparametric approach and one parametric approach) for implementing the MPSE decomposition. After providing a short summary of the approaches, I detail a semiparametric estimation approach in the main text, and refer readers to Supplementary Material A for further detail on the parametric approach). I also provide a simulation study comparing the performance of the two estimation strategies in Supplementary Material B.
The first estimation approach is a semiparametric, debiased machine learning (DML) estimator, which uses efficient influence functions and cross-fitting to estimate each component of the monotonic path-specific effect (MPSE) decomposition. The second is a simpler regression-with-residuals (RWR) estimator, which relies on parametric linear models and sequential residualization to translate the $\theta_{k}$ components in Equation 6 directly into regression coefficients that can be read off from these linear models. The RWR approach provides a transparent, low-computational alternative that directly links the decomposition to regression coefficients, making it useful both as a practical estimator in large datasets and as a parametric benchmark against which to compare the DML results. Nevertheless, when $X$ and $\overline{Z}_{K}$ are high-dimensional, the parametric models underlying RWR may be misspecified, which can in turn introduce bias. By contrast, the DML approach is robust to flexible, high-dimensional models for the treatment, mediators, and outcome.
The DML approach is characterized by two components: first, the use of a Neyman orthogonal estimating equation based on the efficient influence function (EIF) for the target parameters, which makes estimates of the parameter “locally robust" to estimates of the nuisance functions; second, the use of a $K$-fold cross-fitting algorithm (Chernozhukov et al., 2017).